OpenStudy (anonymous):

solve each system by substitution be sure to give your solution as an ordered pair x+y=5 x-y=3

OpenStudy (anonymous):

2x = 8 x =4 4 + y = 5 y = 1 are you still struggling on this topic or checking answers?

OpenStudy (anonymous):

i am checking an answer for this one but i can get confused

OpenStudy (anonymous):

can you help me out with this one x+2y=9 x+y=28

OpenStudy (anonymous):

your object is to remove one of the variables. so you have to find the simpliest way to do that. here i would multiply 2nd by negative 1 x +2y = 9 -x -y = -28 y = -19 plug that back into one either one x + 2y = 9 x + 2(-19) =9 x -38 = 9 x =9 +38 x =47

OpenStudy (anonymous):

how about this question 7x+3y=35 x-2y=5

OpenStudy (anonymous):

ok - you gotta try what do you think we should do?

OpenStudy (anonymous):

times 7 *x to make 8x

OpenStudy (anonymous):

u mean multiply 2nd equation by 7 : ok but lets try -7

OpenStudy (anonymous):

multiply each term in 2nd equation by -7

OpenStudy (anonymous):

okay what did u get im more visual

OpenStudy (anonymous):

cause i am stuck

OpenStudy (anonymous):

i got something weird i think i messed up i got a decimal

OpenStudy (anonymous):